OpenStudy (anonymous):

find the length of the curve x=t^3, y=3t^2/2, 0

OpenStudy (anonymous):

Formula for length of curve is: $$\int_a^b \sqrt{1 + (\frac{\mathrm{d}y}{\mathrm{d}x})^2} \, \mathrm{d}x$$ That won't format correctly, but basically: Integral from a to b of sqrt(1 + (dy/dx)^2) with respect to x.

OpenStudy (anonymous):

hmm... this is for a parametrized curve.. are you sure that formula will work?

OpenStudy (anonymous):

No, it will not work, but it gives us something to build on. Since I'm tired of my notation not showing up properly, read this: http://tutorial.math.lamar.edu/Classes/CalcII/ParaArcLength.aspx

OpenStudy (anonymous):

ahh wait.. i see my error. I calculated dy/dt wrong... integrated instead of taking derivative.

OpenStudy (anonymous):

Aw man, eashy, spoilin' my fun.

OpenStudy (anonymous):

i used L=int(0,sqrt(3))sqrt( (dx/dx)^2+(dy/dt)) dt

OpenStudy (anonymous):

its 7?

OpenStudy (anonymous):

Yes, indeed it is.

OpenStudy (anonymous):

woo thanks! now to move on to polar coordinates .

OpenStudy (anonymous):

That is still one irritating integral.

OpenStudy (anonymous):

i agree. at least this class is easier than calc 2

OpenStudy (anonymous):

At the end of the day, it's all really fun.